# FaceLift: Single Image to 3D Head with View Generation and GS-LRM

### v1

@@ B.2 Additional Results on In-the-wild Images
We present additional results on in-the-wild images in Fig. 14. FaceLift demonstrates the ability to effectively handle diverse hairstyles and beards. Notably, it excels at hallucinating unobservable hairline splits and synthesizing the transparent properties of hair using Gaussians with low opacity. FaceLift reconstructs photo-realistic 3D heads under various lighting conditions and can be further extended to the reconstruction of cartoon characters.
Figure 14: Results of FaceLift on in-the-wild images. FaceLift is capable of reconstructing complex facial hair, including a variety of hairstyles and beards. It can also accurately handle dramatic facial expressions and varying lighting conditions. Additionally, it is well-suited for the reconstruction of cartoon characters.
Appendix C Experimental Details
C.1 Details on Benchmark Evaluation
Test Camera Extrinsic. Both the Cafca [6] and Ava-256 [34] datasets offer multi-view RGB images along with corresponding camera poses. However, their camera systems differ from those used in FaceLift and the baselines. Directly applying their camera poses for inference is infeasible. Hence, we recalculate the test camera extrinsic in each method’s camera system with the following procedure.
@@ C.1 Details on Benchmark Evaluation
The Ava-256 dataset uses a world coordinate system with the origin set at one of the camera positions. We first re-center the world coordinate origin to the midpoint of all camera locations, which is approximately the center of the human head. This step is unnecessary for the Cafca dataset, as its world coordinate origin is defined as the head’s center. Next, we compute the rotation transformation from the test camera pose to the input camera pose within the dataset’s coordinate system. We then apply the same transformation to the input camera pose in each method’s camera system and rescale the translation to match the settings of each method to get the test camera extrinsic under each method’s camera system. After applying the camera pose transformation, perfect alignment is not achieved due to differences in camera distance and intrinsic parameters. To address this, we manually crop and scale the rendered images for closer alignment with the target images.
@@ C.2 Implementation Details
GS-LRM. For fine-tuning GS-LRM, we mostly maintain the original setup described in [64]. During each training step, we randomly sample a set of 8 images (4 as input views and 4 as supervision views) from 32 random HDR environment light renderings. Both input and output images are rendered at a resolution of 512$\times$512. The model is fine-tuned for 20,000 steps using eight A100 GPUs, each equipped with 40 GB of memory.
C.3 Datasets
Cafca Dataset. The Cafca dataset [6] comprises 1,500 identities, 30 camera poses, 13 expressions, and three environments. From this, we select 40 identities, as detailed in Tab. 3. We utilize the first expression and the first environment (folder 00000_000) for each identity. The input view corresponding to each identity is also specified in Tab. 3. In experiments without facial landmark alignment, all 30 camera poses are used. For experiments with facial landmark alignment, the test views employed are outlined in Tab. 3.
@@ C.3 Datasets
Table 3: Identities and views used for the experiment on Cafca dataset [6].
Ava-256 Dataset. The Ava-256 dataset [34] consists of 256 identities, each captured by 80 cameras, with over 5,000 frames per camera. For qualitative evaluation, we select 10 identities, each with 10 test camera views. All selected frames feature natural expressions. We use camera 401168 as the input view, as it captures the front view of the faces and is positioned at the center of Ava-256’s world coordinate system. The input view, test views, and corresponding frame IDs are detailed in Tab. 4.
@@
Table 4: Identities and views used for the experiments on Ava-256 dataset [34].
Appendix D Limitations
FaceLift achieves high-fidelity, photorealistic 3D head reconstruction from a single input image. It provides detailed representations of hair and skin texture while demonstrating superior identity preservation compared to existing methods.
Despite these appealing results, our approach has certain limitations. First, our synthetic dataset does not include accessories such as hats or glasses. As a result, when the input image features a hat, the model may generate hair-like textures to approximate the back of the hat, as illustrated in Fig. 15, row 1. This limitation could be addressed by incorporating synthetic data with accessories. Additionally, when handling out-of-distribution inputs, such as those in Fig. 15, row 2, the model occasionally generates extraneous hair. This issue might be mitigated by refining the training data distribution or introducing text prompts to enhance control over the multi-view diffusion generation process.

### v2

@@ 3.3 Additional Results on In-the-wild Images
We present additional results on in-the-wild images in Fig. 24, Fig. 25 and Fig. 26. FaceLift demonstrates the ability to effectively handle diverse hairstyles and beards. Notably, it excels at hallucinating unobservable hairline splits and synthesizing the transparent properties of hair using Gaussians with low opacity. Our method reconstructs photo-realistic 3D heads under various lighting conditions and can be further extended to the reconstruction of cartoon characters.
3.4 Additional Ablation Study
Figure 16: Training images used in the study of input view reconstruction. We show example images for training baselines w/o. Input View Reconstruction and w. Input View Reconstruction. The difference lies in the elevation of the first target image.
Figure 17: Importance of input view reconstruction. The diffusion model without input view reconstruction training suffers from identity loss. Additionally, it fails to generate accurate face paint (row 1), diverse hair colors (row 2), varied expressions (row 3 and 4), and accessories (row 5).
More Results on Input View Reconstruction.
Baseline PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DreamSim $\downarrow$ ArcFace $\downarrow$
w/o Input View Reconstruction 16.02 0.7884 0.2893 0.1438 0.2367
w/o Multi-view Attention 16.29 0.7885 0.2861 0.1552 0.2126
Full Model 16.61 0.7968 0.2694 0.1096 0.1573
Table 4: Quantitative results of ablation studies. FaceLift achieves better quantitative results with more suitable representations and specialized training data.
We show training samples for two baselines w/o. Input View Reconstruction and w. Input View Reconstruction in Fig. 16. As the target views are different, baseline w/o. Input View Reconstruction is trained to generate six images with novel camera poses, while baseline w. Input View Reconstruction reconstruct the input image and generate five images with novel poses. Inference results on real world images are displayed in Fig. 17 to illustrate the importance of reconstructing the input image during multi-view diffusion training. The results demonstrate that input view regeneration prevents the model from being confined to the training data distribution, thereby enhancing its ability to preserve identity. Quantitative results of baseline w/o. Input View Reconstruction is shown in Tab. 4.
Figure 18: Results of directly applying FaceLift to video input. By processing each video frame independently, FaceLift generates a sequence of Gaussians that preserves consistent visual identity and accurately captures facial expressions. However, this baseline does not consider temporal consistency.
Figure 19: Autoregressive Generation for 4D Rendering. ”AR Gaussians” denotes autoregressively generated Gaussians. With FaceLift, each video frame is independently converted into a 3D Gaussian representation. An anchor frame at timestamp $t$ (highlighted by the blue box) produces Canonical Gaussians $G_{t}$, which are then deformed into the representations for subsequent frames, $G^{\prime}_{t+1}$, $G^{\prime}_{t+2}$, …, etc. This deformation is supervised by the rendered output Gaussians $G_{t+1}$, $G_{t+2}$, …, etc., produced by FaceLift. Iteratively applying this process yields a temporally consistent Gaussian sequence that supports rendering from any viewpoint.
Figure 20: Deformation Network. The deformation network $D_{t}$ is an eight-layer MLP that predicts geometric deformations, including positional shifts, opacity adjustments, and scale changes. Combined with the Gaussian representations from the previous frame $G_{t}^{\prime}$, it forms the Gaussian representation for the next frame $G_{t+1}^{\prime}$.
Figure 21: Results of applying FaceLift on video. Our proposed autoregressive generation pipeline enables FaceLift to be applied directly to video sequences, achieving 4D novel view synthesis – rendering at any given timestamp and camera pose. Video results are shown in the supplementary material.
3.5 Applying FaceLift on Videos
FaceLift can be directly applied to video frames and achieve high-quality facial reconstructions with consistent visual identity and accurate facial expression, as shown in Fig. 18. However, since FaceLift is not trained on video data, many full-head details are generated independently by the diffusion models, resulting in subtle flickering. In this supplemental document, we introduce a simple yet effective method that leverages FaceLift and autoregressive training to achieve high-quality, temporally smooth 4D facial reconstructions.
Given an input video $\{F_{0},F_{1},\dots F_{T}\}$, we process each video frame $F_{t}$ sequentially to generate a set of 3D Gaussian sequences $\{G_{0},G_{1},\dots G_{T}\}$, where each $G_{t}$ represents the obtained Gaussian representation at timestamp $t$. As each $G_{t}$ is generated from frame $I_{t}$ without interaction with other frames, directly rendering from this Gaussian sequence creates artifacts resulting from time-inconsistency. Hence, we propose an autoregressive generation pipeline, as shown in Fig. 19.
We first select an anchor frame at timestamp $t$ (marked with blue box), and treat its corresponding 3D Gaussian splats as the canonical Gaussians $G_{t}$ (marked with blue box). Then, for a following timestamp $t+1$, we train a deformation network $D_{t}$ to predict Gaussian splats $G_{t+1}^{\prime}$ deformed from $G_{t}$ supervised by rendering results from $G_{t}$. The deformation network is an 8-layer MLP, which takes the $x,y,z$ position of each Gaussian in $G_{t}$ as input and predicts $\Delta x$, $\Delta y$, $\Delta z$, opacity change $\Delta\alpha$ and scale change $\Delta s$. These deformation parameters are combined with $G_{t}$ to generate $G_{t+1}^{\prime}$, as shown in Fig. 20.
To train the deformation network, we render six views with the same camera poses as the multi-view diffusion outputs from $G_{t+1}^{\prime}$, and the renderings of the same camera poses from $G_{t+1}$ are used as pseudo ground truth supervision. Then we treat $G_{t+1}^{\prime}$ as the initial Gaussians and train deformation network $D_{t+1}$ to generate $G_{t+2}^{\prime}$. Iteratively, we will get a Gaussian sequence $\{G_{0}^{\prime},G_{1}^{\prime},\dots,G_{T}^{\prime}\}$. Given any timestamp, we can select the corresponding 3D Gaussians from this Gaussian sequence and render from any given pose. The results of this method are shown in Fig. 21, which demonstrate improved temporal consistency while preserving identity and achieving accurate expression modeling. Please refer to the supplementary video for additional video rendering results.
4 Method Details
4.1 Details on View Generation
Given a single near frontal view face image with azimuth $\alpha$, the multi-view diffusion model will generate six views with azimuths equal to $\alpha$, $\alpha\pm 45^{\circ}$, $\alpha\pm 90^{\circ}$, $\alpha+180^{\circ}$}, covering 360 degrees of the human head. All images, both input and generated output, maintain a zero elevation angle, ensuring consistent horizontal viewpoints. The generated views consist of: a reconstructed front view matching the input image; left and right profiles capturing the sides of the head; and a back view that synthesizes hair structure and color based on the frontal input and learned priors. We also generate three-quarter views (left-front and right-front) to enhance facial details in the following reconstruction stage.
To generate unseen views of the human head, we reformulate view synthesis from a single image as a conditional diffusion process. Specifically, we employ a DDPM-based diffusion model $f_{D}$ to generate $N$ distinct views, denoted $X_{0}^{1},X_{0}^{2},\dots,X_{0}^{N}$, from a single front-facing image $y$ and corresponding text embeddings $e^{1},e^{2},\dots,e^{N}$. This process can be expressed as:
$\{X_{0}^{1},X_{0}^{2},\dots,X_{0}^{N}\}=f_{D}\Bigl{(}y,\{e^{1},e^{2},\dots,e^{N}\}\Bigr{)}.$ (5)
Our objective is to learn the joint distribution of these views conditioned on the input image and text embeddings. We denote this joint distribution as:
$p_{\theta}(x_{0}^{1:N}\mid y,e^{1:N})\equiv p_{\theta}\Bigl{(}\{x_{0}^{1},\dots,x_{0}^{N}\}\mid y,\{e^{1},\dots,e^{N}\}\Bigr{)}.$ (6)
In the following discussion, we omit the condition $y$ and $e^{1},e^{2},\dots,e^{N}$ for simplicity. The joint distribution as $p_{\theta}(x_{0}^{1:N})$ is characterized by a Markov Chain (reverse process):
$p_{\theta}(x_{0:T}^{1:N})$ $=p_{\theta}(x_{T}^{1:N})\prod_{t=1}^{T}p_{\theta}\bigl{(}x_{t-1}^{1:N}\mid x_{t}^{1:N}\bigr{)}$ (7)
$=p_{\theta}(x_{T}^{1:N})\prod_{t=1}^{T}\prod_{n=1}^{N}p_{\theta}\bigl{(}x_{t-1}^{n}\mid x_{t}^{1:N}\bigr{)},$
where $p_{\theta}(x_{T}^{1:N})=\mathcal{N}(x_{T}^{1:N};\textsc{0},\textsc{I})$ and $p_{\theta}(x_{t-1}^{n}\mid x_{t}^{1:N})=N(x_{T}^{n};\mu_{\theta}^{n}(x_{t}^{1:N},t),\sigma_{t}^{2}\textsc{I})$. $\mu_{\theta}(x_{t}^{1:N},t)$ is a trainable component while the variance $\sigma_{t}^{2}$ is untrained time-dependent constants. To learn $\mu_{\theta}$ for generation, a Markov chain called forward process is constructed as:
$q\bigl{(}x_{1:T}^{1:N}\mid x_{0}^{1:N}\bigr{)}$ $=\prod_{t=1}^{T}q\bigl{(}x_{t}^{1:N}\mid x_{t-1}^{1:N}\bigr{)}$ (8)
$=\prod_{t=1}^{T}\prod_{n=1}^{N}q\bigl{(}x_{t}^{n}\mid x_{t-1}^{n}\bigr{)},$
where $q\bigl{(}x_{t}^{n}\mid x_{t-1}^{n}\bigr{)}\;=\;\mathcal{N}\bigl{(}x_{t}^{n};\sqrt{1-\beta_{t}}\,x_{t-1}^{n},\,\beta_{t}\mathbf{I}\bigr{)},$ and $\beta_{t}$ are constants. DDPM [22] shows that by defining
$\mu_{\theta}^{n}(x_{t}^{1:N},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}^{n}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\epsilon_{\theta}\bigl{(}x_{t}^{1:N},t\bigr{)}\right).$ (9)
$\alpha_{t}$ and $\bar{\alpha}_{t}$ are constants derived from $\beta_{t}$ and $\epsilon_{\theta}$ is a noise predictor. We learn $\epsilon_{\theta}$ by
$\ell\;=\;\mathbb{E}_{t,x_{0}^{1:N},n,\epsilon^{1:N}}\bigl{[}\|\epsilon^{n}\;-\;\epsilon_{\theta}^{n}(x_{t}^{1:N},\,t)\|_{2}\bigr{]},$ (10)
where $\epsilon^{1:N}$ is the Gaussian noise of size $N\!\times\!H\!\times\!W$ added to all $N$ views, and $\epsilon_{\theta}^{n}$ is the noise predictor on the $n_{th}$ view. We provide ablation study results of the multi-view attention mechanism in Tab. 4.
5 Experimental Details
5.1 Details on Benchmark Evaluation
@@
Test Camera Extrinsic. Both the Cafca [8] and Ava-256 [39] datasets offer multi-view RGB images along with corresponding camera poses. However, their camera systems differ from those used in FaceLift and the baselines. Directly applying their camera poses for inference is infeasible. Hence, we recalculate the test camera extrinsic in each method’s camera system with the following procedure. The Ava-256 dataset uses a world coordinate system with the origin set at one of the camera positions. We first re-center the world coordinate origin to the midpoint of all camera locations, which is approximately the center of the human head. This step is unnecessary for the Cafca dataset, as its world coordinate origin is defined as the head’s center. Next, we compute the rotation transformation from the test camera pose to the input camera pose within the dataset’s coordinate system. We then apply the same transformation to the input camera pose in each method’s camera system and rescale the translation to match the settings of each method to get the test camera extrinsic under each method’s camera system. After applying the camera pose transformation, perfect alignment is not achieved due to differences in camera distance and intrinsic parameters. To address this, we manually crop and scale the rendered images for closer alignment with the target images.
@@ 5.2 Implementation Details
Transformer-based Gaussian Reconstructor. The training of the reconstructor follows [74]. During each training step, we randomly sample a set of 8 images (4 as input views and 4 as supervision views) from either 32 ambient light renderings or 25 random HDR environment light renderings. Both input and output images are rendered at a resolution of 512×512. The model is fine-tuned for 20,000 steps using eight A100 GPUs, each equipped with 40 GB of memory.
Figure 22: Visual Comparison with LGM. Leveraging the outputs of our multi-view diffusion model enhances the performance of LGM [56] (denoted as Our MV + LGM). We further fine-tuned LGM using our synthetic human head data, resulting in Our MV + Fine-tuned LGM; however, its performance was inferior to that achieved with the original weights in Our MV + LGM.
For a fair comparison, we also fine-tune LGM [56] with our synthetic data with their provided training codes. However, the fine-tuned LGM achieves inferior performance than the original weights, as shown in Fig. 22.
5.3 Datasets
Cafca Dataset. The Cafca dataset [8] comprises 1,500 identities, 30 camera poses, 13 expressions, and three environments. From this, we select 40 identities, as detailed in Tab. 5. We utilize the first expression and the first environment (folder 00000_000) for each identity. The input view and test views corresponding to each identity are also specified in Tab. 5.
Ava-256 Dataset. The Ava-256 dataset [39] consists of 256 identities, each captured by 80 cameras, with over 5,000 frames per camera. For qualitative evaluation, we select 10 identities, each with 10 test camera views. All selected frames feature natural expressions. We use camera 401168 as the input view, as it captures the front view of the faces and is positioned at the center of Ava-256’s world coordinate system. The input view, test view, and corresponding frame IDs are detailed in Tab. 6.
6 Limitations
FaceLift achieves high-fidelity, photorealistic 3D head reconstruction from a single input image. It provides detailed representations of hair and skin texture while demonstrating superior identity preservation compared to existing methods. Despite these appealing results, our approach has certain limitations. First, our synthetic dataset does not include accessories such as hats or glasses. As a result, when the input image features a hat, the model may generate hair-like textures to approximate the back of the hat, as illustrated in Fig. 23, row 1. This limitation could be addressed by incorporating synthetic data with accessories. Additionally, when handling out-of-distribution inputs, such as those in Fig. 23, row 2, the model occasionally generates extraneous hair. This issue might be mitigated by refining the training data distribution or introducing text prompts to enhance control over the multi-view diffusion generation process. Finally, in some cases, the unseen regions of the face appear more blurred than the visible areas (frontal face). Our system emphasizes detailed reconstruction of the front face: most views generated by the diffusion model concentrate on the frontal region, and the input-view reconstruction strategy faithfully preserves its features. In contrast, features of the back of the head are primarily learned from synthetic data. Additionally, when simulating lighting, the model tends to darken the back head and introduce shadows, often causing the hair to appear black.
Figure 24: Results of FaceLift on in-the-wild images. FaceLift excels at reconstructing intricate and diverse facial hair, encompassing a wide array of hairstyles and hair colors. It also accurately captures a broad range of skin tones.
Figure 25: Results of FaceLift on in-the-wild images. FaceLift also demonstrates the ability to reconstruct faces exhibiting a wide range of pose variations. It can also accurately handle extreme expressions.
Figure 26: Results of FaceLift on in-the-wild images. FaceLift realistically reconstructs detailed facial textures. Additionally, FaceLift is well-suited for reconstructing cartoon characters.
@@ 6 Limitations
Table 5: Identities and views used for the experiment on Cafca.
@@
Table 6: Identities and views used for the experiments on Ava-256.